TMS Meeting Minutes, June 10, 2021

Attendees: Andrew, Breann, Heejoo, John, Matthieu, Olga, Trenton, Yang (Zoom Meeting)

Summer Shutdown
  • Early state of summer shutdown
  • July 12th is the start of summer shutdown
  • We will have to make a trip before that or just after
  • Summer shutdown ends on August 31st, so we need to reinstall everything before that
  • We will need to make a plan for the TMS re-install
  • This time, we won't be using multi-core fiber, as it has some issues, instead just single core
  • Will replace broken fibers, and some weak channels (some channels have <5% return)
  • 4-5 channels have broken fibers right now
  • Will also conduct laser tracker survey
TMS Testing (Engineering Night, Passive Mode)
  • Olga conducted offline FPIA, and summarized data
  • Heejoo to do some post-processing of the data to correlate with mirror temps, etc.
  • Planning to provide data by the end of this week or early next week

  • Thanks to Olga for sending the notice to observers that we plan to use Olga during the observing night
  • If you have any questions or want to see any data from the past week, reach out to Heejoo
  • Andrew would like to see the channel failure statistics (from the std. dev. filter, etc.)
  • Andrew suspects that the bimodal click tuning of the collimator mounts may be the cause of some channel dropping
  • Proposes using a small tack of super glue to lock the mounts to the appropriate range of rotation
  • Please don't put any super glue on the primary mirror
  • John Hill suggests using gluptol (not sure how that's spelled) to provide the necessary constraint
  • Andrew notes that this is a short term solution, and that there may be need for a redesign of the collimator mounts
  • Long term stability directly impacts channel stability and drops
  • Heejoo notes that there are three channels which were consistently low return, but they appear to have improved after some maintenance
  • Andrew suggests including a channel health indicator to the data output formatting, giving a percentage of failure for each channel throughout the night


  • Noticed a new TMS error message
  • In the past, our errors have been caused by sending too many commands in too short a period of time
  • This time, the laser tuning speed could not be set
  • The TMS machine left the shutter open for a full day until Heejoo clicked okay on all the error messages
  • It seems that if there is an issue during communication or during the loop, there is no way to close the shutter itself
  • Waiting for a response from Etalon about how they could implement this
  • As it stands now, the laser meter does not send an alert when the laser shutter has been open for a long time, so there is no alert when the system fails and the shutter stays open in error
  • There should never be a case when the shutter is just constantly open, so we need some sort of alert for a continous on mode of operation
  • Something like, the last ten/three sets of measurements show always on, red flag
  • The current TMS laser warning system has a range threshold, such as voltage less than 0.2 is off
  • If the voltage is greater than this but less than the on case, the power meter is turned off
  • If above the on voltage, reports that the laser is on
  • Matthieu requests that we send him whatever it is we need from iMonnit
  • Also a suggestion from Andrew to consider the possibility of purchasing a new device with higher sampling may be worthwhile
  • Matthieu's team to look into providing a better solution for the laser power meter
  • John suggests a simple algorithm to send out another alert if we get three on readings in a row

06/07 data was interrupted by laser state error

06/09 Passive Data
  • Needs to be found
  • LBT TMS Commissioning Plan
  • See most recent document 2021.06.10... for updates
  • Please provide your contributions
Kill-a-watt Measurements
  • Measurements were made
  • Came out to about 0.45 kWh average
  • Will be developing a plan for the summer shutdown hardware

  • Tonight, there might be an opportunity for TMS use
  • TMS will only be used if the conditions are not good enough for the other projects
  • The outstanding gap in our on-sky dat is having continuous Zernikes from on-sky data
  • Would be very useful to have FPIA run, take a TMS reference, then go to another field and take a reference
  • Oscillating between two fields could produce very interesting results, and could be helpful for verifying performance
  • Olga wonders what sort of conditions are required for these tests
  • Andrew comments that, if we are in 3 arcsec seeing, there isn't much point in taking Zernike's
  • However, 1.5 arcsec seeing could give a useful idea of Zernikes
  • Getting hours of Zernike data while actively running TMS is high priority right now
  • After converging with FPIA, impart an offset of 800 microns in Z, take reference, slew to in-focus OB, and then run TMS to hold the offset while taking pupils

  • Olga is putting together a document for observers regarding TMS usage
  • Will be sending the document link to the TMS mailing list
  • Andrew just found another error on the Multiline PC...
  • Heejoo about to restart the Multiline system to wipe any errors that remain

-- TrentonBrendel - 11 Jun 2021
Topic revision: r2 - 16 Sep 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback